The solutions to a quadratic equation are -4 and 4. What is the equation of its axis of symmetry?

Answer by lwsshak3(11628)

The solutions to a quadratic equation are -4 and 4. What is the equation of its axis of symmetry?
**
Solutions to quadratic equations are zeros or x-intercepts
Given equation: y=(x+4)(x-4)=x^2-16 (difference of 2-squares)
Since this is a parabola that opens upwards, equation of axis of symmetry is x=0 (vertical line between -4 and 4 on the x-axis)